OpenStudy (anonymous):

a rectangular yard is 20ft by 15ft. the yard is covered with grass except for an 8.5ft square flower garden. how much grass is in the yard? answer: 227.75i dont get how they got that answer

OpenStudy (anonymous):

 This is what you're given. To find how much grass there is, you find the area of the rectangular yard and the area of the flower garden. Area = length *width A = 20*15 What is 20*15, @kelewood? Next, you find the area of the square. Area = side^2 A = 8.5^2 What is 8.5^2? Once you've found those numbers, subtract the area of the flower garden from the rectangular yard.